OpenStudy (eric_d):

sketch following curve,showing the asymtotes a)y=3^x-1 @ikram002p

OpenStudy (ikram002p):

start by sketch y=3^x can u do that ?

OpenStudy (eric_d):

how

OpenStudy (ikram002p):



OpenStudy (eric_d):

ok is it because u put x=0

OpenStudy (eric_d):

I mean u put x=0 so y=1 3^0=1 am I rite

OpenStudy (ikram002p):

yes :D

OpenStudy (eric_d):

ok..what's next

OpenStudy (ikram002p):

next u have y=3^x-1 so only shift ur skitch down one unit like this